PROSEC Find all relative extreme points and saddle points of f(x, y) = x - y2 - 2xy.
Answer #1

. Given that flny) = y ² any Here First of all we have to find all first and and order partial derivative Fu= 3 ay B and funFor value of y then When n o y=0 and when u=-2 then ya g it rom eq @ 29=- 24 39 u =-y . y=-u - so critical points are (0,0) aNow let we checkin D(0,0) = 0x-2-(-2) ? = cuco so point 1010) is saddle point nrow at point la 01-31, ) = Fuel , ). Fay ( *)